OpenStudy (anonymous):

hey can anyone help me with series/sequences? (algebra 2)

OpenStudy (sandra):

hey what type of sequences/series?

OpenStudy (anonymous):

um not sure can i give u a problem and u try to help?

OpenStudy (anonymous):

????

OpenStudy (sandra):

sure

OpenStudy (anonymous):

Write the series using summation notation. 7 + 10 + 13 + 16 + 19

OpenStudy (sandra):

ah ok. so the first number is 7, so that is a constant

OpenStudy (anonymous):

k...

OpenStudy (anonymous):

do i have to pay to get help on here??

OpenStudy (sandra):

\[\sum_{n=0}^{n=4} 7+3*n\] so something liek that?

OpenStudy (anonymous):

yes but how did u get that....

OpenStudy (sandra):

well, you are trying to make the summation equation part of it (the 7 + 3*n in this case), generate all of the numbers in the series if you input different values for "n". So given 7+10+13+16+19, you can start at the left. So if when n=0, the thing on the right is 7, then we know that 7 is a part of the summaition series

OpenStudy (anonymous):

oh

OpenStudy (sandra):

so then the question is , what can I add to 7, so when n=1 I'm adding 10, when n=2 I'm adding 13, and n=3 I'm added 16

OpenStudy (sandra):

so in fact my answer was wrong =)

OpenStudy (anonymous):

o...

OpenStudy (sandra):

or wait, sorry, I was right lol

OpenStudy (sandra):

ok so if n=0, and the right side is 7

OpenStudy (sandra):

and n=1, and the right side is 10

OpenStudy (sandra):

I've added 3

OpenStudy (sandra):

so my guess is that the right side is 7+3(n)

OpenStudy (sandra):

let's see if it holds true for n=2

OpenStudy (sandra):

for n=2, according to me, the third number is the series would be: 7 + (3*2) = 7+6 = 13

OpenStudy (sandra):

and so on..

OpenStudy (anonymous):

kk thanks!!!!!!!

OpenStudy (anonymous):

do i have to pay for this..

OpenStudy (anonymous):

no its free